bonez:

What is the average atomic mass of the element? An element, X, has two isotopes, X-14 and X-15. Use the data in the table to find the average atomic mass of element X. Isotope Atomic Mass Abundance (%) X-14 14.003 99.636 X-15 15.000 0.364 The average atomic mass of element X is amu.

ZoeyBiocth16:

Average atomic mass of element X is 14.0066 amu.

ZoeyBiocth16:

average atomic mass = mass of first isotope(abundance) + mass of second isotope(abundance) mass of first isotope is 14.003 and it's abundance is 0.99636. Mass of second isotope is 15.000 and it's abundance is 0.00364. Let's plug in the values in the formula and do calculations: average atomic mass = 14.003(0.99636) + 15.000(0.00364) average atomic mass = 13.9520 + 0.0546 average atomic mass = 14.0066 amu